# spacing=nonfrench

### v1

@@ III-D Throughput
Figure 7: Memory requirement with batching in prompt and token phases.
@@ III-E Memory utilization
During an LLM inference, the GPU memory is used to host the model weights and activations, as well as the KV caches (Section II-B). As the number of tokens in a batch increases, the memory capacity requirements for the KV cache also increase. Figure 7 shows the memory capacity utilization during each phase as the number of tokens in the batch increases. During the prompt phase, the active prompt tokens create the KV cache according to their size. During the output token phase, each active generated token that is being processed accesses the KV cache of its entire context so far.
@@ III-F Power utilization
When hosting machines, cloud providers need to consider the peak power draw, which has a direct impact on the datacenter cost [21]. This is especially important when building clusters for GPUs since they consume much higher power than regular compute machines [37, 36]. Figure 8 shows the GPU power draw normalized to the thermal design power (TDP) when running prompt and token generation phases. As the the prompt phase is compute intensive, the power draw increases with the batch size. On the other hand, the token phase is memory bound and the power draw does not vary when increasing the number of tokens to process.
Providers can cap the power usage of the machines to reduce the peak power. Figure 9 shows the impact on latency when increasing the power caps for both prompt and token phases. The prompt phase is highly sensitive to the power cap and the latency increases substantially. On the other hand, the token generation phase sees almost no impact in latency when power capping by over 50% (i.e., 700 to 350W).
@@
TTFT 185 ms 95 ms 0.51$\times$ 155 ms 84 ms 0.54$\times$
TBT 52 ms 31 ms 0.70$\times$ 40 ms 28 ms 0.70$\times$
E2E 856 ms 493 ms 0.58$\times$ 4957 ms 3387 ms 0.68$\times$
Cost [3] \$0.42 \$0.52 1.24$\times$ \$2.4 \$3.6 1.50$\times$
Energy 1.37 Whr 1.37 Whr 1.00$\times$ 7.9 Whr 9.4 Whr 1.20$\times$
TABLE IV: P50 request metrics on A100 vs. H100 without batching on Llama2-70B.
@@ III-G GPU hardware variations
Given the different characteristics of prompt and token generation phases, we measure the performance impact on the two from running on different hardware. Table I shows the specifications for DGX-A100 [34] and DGX-H100 [14]. The memory to compute ratio favors A100 compared to H100. Table IV shows our findings. We see a lower performance impact on the token generation phase (TBT) as compared to the Prompt phase (TTFT). Since coding requests are dominated by prompt phase, by having very few generated tokens, the E2E latency impact from A100 is worse on coding than conversation. Furthermore, we see that the overall inference cost and energy are better or equal on A100, compared to H100.
Insight VII: The generation phase can run on cheaper, lower maximum performance hardware for better Perf/W and Perf/\$ efficiencies.
@@ IV Splitwise
Based on the insights presented in Section III, we propose Splitwise: a technique to split the prompt and generation phases in the LLM inference on to separate machines.
Figure 10 shows the high-level overview of Splitwise. We maintain two separate pools of machines for prompt and token processing. A third machine pool, the mixed pool, expands and contracts as needed by the workload. All the machines are pre-loaded with the model of choice. When a new inference request arrives, the scheduler allocates it to a pair of machines (i.e., prompt and token). The prompt machines are responsible for generating the first token for an input query, by processing all the input prompt tokens in the Prompt phase and generating the KV-cache. The prompt machine also sends over the KV-cache to the token machine, which continues the token generation until the response is complete. We use continuous batching at the token machines, to maximize their utilization. The mixed pool is the only set of machines where mixed batches would apply. We use mixed continuous batching in those machines.
@@
Splitwise uses a hierarchical two-level scheduling as shown in Figure 10. The cluster-level scheduler (CLS) 1 is responsible for routing the incoming requests to particular machines and re-purposing of machines. The machine-level scheduler (MLS) 2 maintains the pending queue and manages batching of requests at each machine.
@@ IV-A Cluster-level scheduling
Request routing. Each machine communicates to the CLS any change in its memory capacity or pending queue. Note that this does not necessarily happen at every single iteration boundary. Then, CLS uses Join the Shortest Queue (JSQ) scheduling [50, 25] to assign a prompt and a token machine to each request upon arrival. We assign the token machine upon arrival to minimize the KV-cache transfer overhead (Section IV-C).
Machine pool management. The CLS mainly maintains two machine pools: the prompt and token pools. Splitwise initially assigns machines to a pool depending on the input/output token distribution and the expected load (i.e., requests per second). If these values deviate considerable from the initial assumptions, Splitwise employs a coarse granularity re-purposing of machines and moves machines between the prompt and token pools. Re-purposing a machine could imply reloading the model (e.g., different degree of parallelism - TP or PP) which can take a significant overhead (e.g., 5 minutes). Therefore, this is done very infrequently if more than 10% of the machines stay in the mixed pool for a long time threshold. The repurposing is performed one-by-one, and during times of lower utilization on the cluster.
Mixed machine pool. To meet SLOs and avoid any performance cliffs due to fragmentation at higher loads, Splitwise maintains a special mixed pool 3. This pool grows and shrinks dynamically, without any noticeable pool-switching latency, based on the request rates and distributions of input and output tokens. A machine in the mixed pool still retains its original identity as a prompt or token machine and goes back to its original pool once there are no tasks of the opposite kind in its pending queue.
If the CLS tries to assign a prompt and token machine for a request using JSQ and it finds that the queue in the selected machine is beyond a threshold, it looks for target machines in the mixed pool. If the mixed pool is also full, it proceeds to look in the opposite pool (i.e., a token machine to run prompts and vice versa) and moves the machine into the mixed pool. Machines in the mixed pool operate exactly as a non-Splitwise machine would, with mixed batching. Once the queue of mixed requests is drained, CLS transitions the machine back to its original pool. For example, when the queue is too long, we can move a prompt machine to the mixed pool to run tokens; once the machine is done running tokens, we transition the machine back into the prompt pool.
@@ IV-B Machine-level scheduling
The MLS runs on each machine and is responsible for tracking the GPU memory utilization, maintaining the pending queue 4, selecting the batch size and the batched requests for each iteration, and reporting the relevant status to the CLS.
@@
Mixed machines. To meet the SLOs for TTFT, the MLS needs to prioritize running prompts and schedules any new prompts in the pending queue immediately. If the machine was running a token phase and there is no capacity, the MLS will preempt tokens. To avoid starvation of the token phase due to preemption, we increase the priority of the token with age and limit the number of preemptions each token can have.
@@ IV-C KV-cache transfer
As discussed in Section II, the KV-cache is generated during the prompt phase of the request, and constantly grows during its token generation phase. In Splitwise, we need to transfer the KV-cache from the Prompt machine to the Token machine 5 (shown in Figure 10) to avoid any duplicate computation. This is the main overhead that Splitwise adds on the LLM inference cluster. In this section, we discuss the impact of KV-cache transfer and how we optimize it.
@@
The time required for the transfer depends on the size of the KV cache (which is directly proportional to the number of prompt tokens) and on the bandwidth of the interconnect between the prompt and the token machines. Even when using fast InfiniBand links between the machines, the overhead for large prompt sizes can easily be a significant fraction of the TBT. We discuss this further in the evaluation in Section V-B.
In Splitwise, we optimize the KV-cache transfer by overlapping it with the computation in the prompt phase. As each layer in the LLM gets calculated in the prompt machine, the KV cache corresponding to that layer is also generated. At the end of each layer, we trigger an asynchronous transfer of the KV-cache for that layer while the prompt computation continues to the next layer. Figure 10(b) shows this asynchronous transfer which reduces the overheads. Layer-wise transfers also allow other optimizations, such as earlier start of the token phase in the token machines, as well as earlier release of KV-cache memory on the prompt machines.
For small prompt sizes, the KV-cache is small and it is unnecessary to pay the overheads of fine-grained layer-wise synchronization required by per-layer transfer. Since the number of tokens in a batch is already known at the start of computation, Splitwise picks the best technique for KV-cache transfer. For small prompts, Splitwise uses serialized KV-cache transfer. On the other hand, for larger prompts sizes, it uses the per-layer transfer.
@@
Splitwise-AA DGX-A100 1.00$\times$ 1.00$\times$ DGX-A100 1.00$\times$ 1.00$\times$ 1$\times$
Splitwise-HH DGX-H100 2.35$\times$ 1.75$\times$ DGX-H100 2.5$\times$ 1.75$\times$ 2$\times$
Splitwise-HHcap DGX-H100 2.35$\times$ 1.75$\times$ DGX-H100 2.5$\times$ 1.23$\times$ 2$\times$
Splitwise-HA DGX-H100 2.35$\times$ 1.75$\times$ DGX-A100 1.00$\times$ 1.00$\times$ 1$\times$
@@ IV-D Provisioning with Splitwise
Type of machines. We primarily propose four variants of Splitwise-based systems: Splitwise-HH, Splitwise-HHcap, Splitwise-HA, and Splitwise-AA. The nomenclature is simply drawn from the first letter representing the Prompt machine type, and the second letter representing the Token machine type. “H” represents a DGX-H100 machine, “A” represents a DGX-A100 machine, and “Hcap” represents a power-capped DGX-H100 machine. Table V shows a summary of the cost, power, and hardware in each of our evaluated systems.
@@
Splitwise-HH uses DGX-H100 for both prompt and token machine pools while Splitwise-AA uses DGX-A100 for both pools. These two variants represent the commonly available setups in providers where machines are homogeneous and interchangeable.
Splitwise-HHcap uses DGX-H100 machines for both prompt and token machines. However, we power cap the token machines down to 70% of their rated power, with each GPU capped by 50% of the power. We propose this design based on Figure 9 and Insight VII (i.e., the prompts phase is impacted by power caps while token has no performance impact with 50% lower power cap per GPU).
@@
Splitwise-HA uses DGX-H100 type for prompt machines and DGX-A100 for the token pool. We choose this configuration based on Table IV, and the Insight VII (i.e., A100s can be more cost- and power-efficient for the token phase).
Number of machines. The LLM inference deployment cluster will need to be sized appropriately with the right number of Prompt and Token machines.
@@
Our methodology involves searching the design space using our event-driven cluster simulator (described in detail in Section V-A). We need to provide (1) the target cluster design (e.g., Splitwise-HA or Splitwise-HHcap), (2) the performance model based on the TTFB and TBT latency for various batch sizes for the specific LLM, (3) a short trace derived from the target prompt and token size distributions for the service (e.g., Figure 3), (4) the SLOs (e.g., Table VI), (5) the constraints (e.g., throughput), and (6) the optimization goal (e.g., minimize cost). Using this information, our provisioning framework searches the space for the desired optimal point. For instance this search with a constraint on maximum achievable throughput and an optimization goal of minimizing cost gives us iso-throughput cost-optimized cluster provisioning across different designs.

### v2

@@ III-D Throughput
Figure 7: Required memory with batching in prompt/token phases.
@@ III-E Memory utilization
During an LLM inference, the GPU memory is used to host the model weights and activations, as well as the KV caches (Section II-B). As the number of tokens in a batch increase, the memory capacity required for the KV cache also increases. Figure 7 shows the memory capacity utilization during each phase as the number of tokens in the batch increases. During the prompt phase, the input prompt tokens generate the KV cache. During the output token phase, each active generated token that is being processed accesses the KV cache of its entire context so far.
@@ III-F Power utilization
When hosting machines, cloud providers need to consider the peak power draw, which has direct impact in the datacenter cost [26]. This is especially important when building GPU clusters, since GPUs consume much higher power than regular compute machines [63, 64]. Figure 8 shows the GPU power draw normalized to the thermal design power (TDP) when running prompt and token generation phases. Since the the prompt phase is compute intensive, its power draw increases with batch size. On the other hand, the token phase is memory bound and its power draw does not vary when increasing the number of tokens to process.
Providers can cap the power usage of the machines to reduce the peak power. Figure 9 shows the impact to latency when increasing the power caps for both prompt and token phases. The prompt phase is highly sensitive to the power cap and the latency increases substantially. On the other hand, the token generation phase incurs almost no latency impact when power capping by over 50% (i.e., 700 to 350W).
@@
[…]
TABLE IV: P50 request metrics on A100 vs. H100 without batching on Llama-70B.
@@ III-G GPU hardware variations
Given the different characteristics of prompt and token generation phases, we measure the performance impact on the two from running on different hardware. Table I shows the specifications for DGX-A100 [15] and DGX-H100 [16]. The memory-to-compute ratio favors A100 over H100. Table IV shows our findings. We see a lower performance impact on the token generation phase (TBT) as compared to the Prompt phase (TTFT). Since coding requests are dominated by prompt phase, by having very few generated tokens, the E2E latency impact from A100 is worse on coding than conversation. Furthermore, we see that A100 has better or equal inference cost and energy overall compared to H100.
Insight VII: Token generation can be run on less compute-capable hardware for better Perf/W and Perf/\$ efficiencies.
@@ IV Splitwise
Based on our characterization insights, we propose Splitwise, a technique to split the prompt and generation phases in the LLM inference on to separate machines.
Figure 10 shows the high-level overview of Splitwise. We maintain two separate pools of machines for prompt and token processing. A third machine pool, the mixed pool, expands and contracts as needed by the workload. All machines are pre-loaded with the model of choice. When a new inference request arrives, the scheduler allocates it to a pair of machines (i.e., prompt and token). The prompt machines are responsible for generating the first token for an input query, by processing all the input prompt tokens in the prompt phase and generating the KV-cache. The prompt machine also sends over the KV-cache to the token machine, which continues the token generation until the response is complete. We use continuous batching at the token machines to maximize their utilization. Machines in mixed pool use mixed continuous batching.
@@
Splitwise uses a hierarchical two-level scheduling as shown in Figure 10. The cluster-level scheduler (CLS) 1 is responsible for machine pool management and for routing incoming inference requests. The machine-level scheduler (MLS) 2 maintains the pending queue and manages batching of requests at each machine.
@@ IV-A Cluster-level scheduling
Machine pool management. The CLS maintains the prompt, token, and mixed machine pools 3. Splitwise initially assigns machines to the prompt or token pool depending on the expected request load and input/output token distributions. Machines from the prompt or token pools may be dynamically moved into and out of the mixed pool to reduce fragmentation and meet SLOs at higher loads. A machine in the mixed pool retains its identity as a prompt or token machine and goes back to its original pool once there are no tasks of the opposite kind in its pending queue. Switching pools does not incur any noticeable latency. If the load distribution deviates considerably from initial assumptions, Splitwise employs a coarse grained re-purposing of machines and moves machines between the prompt and token pools. Re-purposing of machines is done infrequently, typically only if they stay in the mixed pool for a considerable amount of time.
Request routing. CLS uses Join the Shortest Queue (JSQ) scheduling [85, 39] to assign a prompt and a token machine to each request. Queue lengths are defined by the number of pending tokens. Each machine regularly communicates to the CLS changes in its memory capacity or pending queue. Note that this does not happen at every iteration boundary. We simultaneously assign both the prompt and token machine when scheduling requests, since we can then overlap KV-cache transfers with prompt computation to reduce transfer overheads (Section IV-C).
When routing requests, if the pending queue is bigger than a certain threshold, the CLS looks for target machines in the mixed pool. If the mixed pool is also full, it proceeds to look in the opposite pool (i.e., a token machine to run prompts and vice versa) and moves the machine into the mixed pool. Machines in the mixed pool operate exactly as a non-Splitwise machine would, with mixed batching. Once the queue of mixed requests is drained, the CLS transitions the machine back to its original pool. For example, when the queue is too long, we can move a prompt machine to the mixed pool to run tokens; once the machine is done running tokens, we transition the machine back into the prompt pool.
@@ IV-B Machine-level scheduling
The MLS runs on each machine and is responsible for tracking the GPU memory utilization, maintaining the pending queue 4, deciding the batch for each iteration, and reporting the relevant status to the CLS.
@@
Mixed machines. To meet the TTFT SLO, the MLS must prioritize running prompts and schedule any new prompts in the pending queue immediately. If the machine is running token phases and has no additional capacity to run the prompt phase, the MLS will preempt tokens. To avoid starvation of the token phase due to preemption, we increase the priority of the token with age and limit the number of preemptions that each request can have.
@@ IV-C KV-cache transfer
As discussed in Section II, the KV-cache is generated during the prompt phase of the request, and it continuously grows during the token generation phase. In Splitwise, we need to transfer the KV-cache from the prompt machine to the token machine 5 (shown in Figure 10) to complete the inference. This transfer delay is the main overhead associated with Splitwise. In this section, we discuss the impact of KV-cache transfer and how we optimize it.
@@
The time required for the transfer depends on the size of the KV cache (which is directly proportional to the number of prompt tokens) and on the bandwidth of the interconnect between the prompt and the token machines. Even when using fast InfiniBand links, the transfer overhead for large prompt sizes could become a significant fraction of the TBT.
In Splitwise, we optimize the KV-cache transfer by overlapping it with the computation in the prompt phase. As each layer in the LLM gets calculated in the prompt machine, the KV cache corresponding to that layer is also generated. At the end of each layer, we trigger an asynchronous transfer of the KV-cache for that layer while the prompt computation continues to the next layer. Figure 11(b) shows this asynchronous transfer which reduces the transfer overheads. Layer-wise transfer also enables other optimizations, such as earlier start of the token phase in the token machines, as well as earlier release of KV-cache memory on the prompt machines.
Layer-wise KV-cache transfer happens in parallel with the prompt computation for the next layer. This requires fine-grained synchronization per layer for correctness. Thus, it is possible to incur performance interference and increase the TTFT, especially for smaller prompts. However, for small prompts the total KV-cache size is small and does not need the layer-wise transfer to hide the latency. Since the number of tokens in a batch is already known at the start of computation, Splitwise picks the best technique for KV-cache transfer. It uses serialized KV-cache transfer for smaller prompts and layer-wise transfer and for larger prompts. We show that the overall transfer and interference overheads are relatively small in Section VI-A.
@@
[…]
@@ IV-D Provisioning with Splitwise
We leverage Splitwise to optimize LLM inference cluster deployments for power, cost, and throughput.
@@
Type of machines. We propose four main variants of Splitwise-based systems: Splitwise-AA, Splitwise-HH, Splitwise-HA, and Splitwise-HHcap. The nomenclature is simply drawn from the first letter representing the Prompt machine type, and the second letter representing the Token machine type. “A” represents a DGX-A100 machine, “H” represents a DGX-H100 machine, and “Hcap” represents a power-capped DGX-H100 machine. Table V shows a summary of the cost, power, and hardware in each of our evaluated systems.
@@
Splitwise-AA uses DGX-A100 for both prompt and token pools, while Splitwise-HH uses DGX-H100 for both. These two variants represent the commonly available setups in providers where machines are homogeneous and interchangeable.
@@
Splitwise-HA uses DGX-H100 for the prompt pool and DGX-A100 for the token pool. We choose this configuration based on Table IV, and the Insight VII (i.e., A100s can be more cost- and power-efficient for the token phase).
Splitwise-HHcap uses DGX-H100 machines for both prompt and token pools. However, we power cap the token machines down to 70% of their rated power, with each GPU capped by 50% of the power. We propose this design based on Figure 9 and Insight VII (i.e., the prompts phase is impacted by power caps while token has no performance impact with 50% lower power cap per GPU).
@@
Number of machines. The LLM inference cluster deployment must be sized with the appropriate number of prompt and token machines. Our methodology involves searching the design space using our event-driven cluster simulator, which is described in detail in Section V. We need to provide as input: (1) the target cluster design (e.g., Splitwise-HA or Splitwise-HHcap), (2) an LLM-specific performance model that can estimate the TTFT and TBT at various input, output, and batch sizes, (3) a short trace derived from the target prompt and token size distributions for the service (e.g., Figure 3), (4) the SLOs (e.g., Table VI), (5) the constraints (e.g., throughput), and (6) the optimization goal (e.g., minimize cost). Using this information, our provisioning framework searches the space for the desired optimal point. For example, searching with a throughput constraint and a cost minimization goal gives us iso-throughput cost-optimized clusters across different designs.
Figure 12: Design space for provisioning a Splitwise-$H$ cluster. Cluster configurations targets a peak throughput of 70 RPS. The cost-optimal Splitwise-$H$ configuration is marked with $\star$ (27 prompt and 3 token machines).
Search space. Figure 12 shows an example of the two-dimensional search space for the number of prompt and token machines under Splitwise-HH for the coding workload (using a 2-minute trace). The simulator outputs the various percentiles for TTFT, TBT, and E2E latencies. Then, we select the clusters that meet the SLOs for each of these metrics and optimize our target function. For example, Figure 12 shows a $\star$ for the setup with 27 prompt and 3 token machines with the lowest cost that achieves 70 RPS. We call this setup iso-throughput cost-optimized.
Optimization. We can use three optimization goals: throughput, cost, and power. Throughput optimization is important for both, the cloud service provider (CSP) and the user. Cost optimization has different importance levels to the CSP and the user. For the CSP, a higher cost for the same throughput might be acceptable if there are gains in power and space requirements for the cluster. However, for the end-user, a higher cost at the same throughput is generally unacceptable. Finally, power optimization is attractive for a CSP, since it enables more GPUs to be deployed in the same datacenter [62, 63], but it may not be as important to the user. We only consider the provisioned power, and not the dynamic power utilization, in our study.
IV-E Practical Considerations
Accuracy impact. Splitwise does not impact accuracy since it uses lossless KV-cache transfer and does not add any randomization. It executes inference with the same parameters and state as on a single machine.
Scalability. Since LLM requests are much longer than typical ML requests [37, 38], they incur lower scheduling overhead for similar cluster sizes. However, the CLS may become a scalability bottleneck for large clusters. Insights from prior work on partitioned or replicated scheduling could help improve scalability [61, 27, 72] and are orthogonal to Splitwise.
Reliability and fault tolerance. If the prompt or the token machine fail, Splitwise simply restarts requests from scratch, similar to today’s LLM serving systems [51, 44]. Alternatively, Splitwise could checkpoint the KV-cache generated after prompt computation into an in-memory database. To recover, Splitwise can use this cache to skip prompt recomputation, and start right away with the token phase. The KV-cache could also be checkpointed periodically during the token phase. Designing safe and efficient failure recovery is out of scope for our paper.
V Methodology
V-A Experimental setup
To evaluate our proposal on real hardware, we implement Splitwise’s KV-cache transfer mechanism on top of vLLM [51]. Our implementation is open source [1]. We run this modified vLLM on two DGX-A100 and two DGX-H10 virtual machines (VMs) on Microsoft Azure with specifications from Table I. These are the VMs used to collect the characterization data in Section III. These machines are connected with InfiniBand and the DGX-H100s have double the bandwidth (i.e., 400 Gbps).
Our implementation of the Splitwise technique assigns machines either a prompt role, or a token role. As the prompt machine generates the first token, it transfers the KV-cache to the token machine using the technique described in Section IV-C. We use MSCCL++ [11], an optimized GPU-driven communication library, to implement the naive and layer-wise KV cache transfers.
In our implementation, the prompt machine uses the zero-copy one-sided put primitive of MSCCL++ to send KV-cache data over InfiniBand as soon as it is ready, without requiring the token machine to issue any receive instructions. Once we have issued a put for all layers, the prompt machine signals a semaphore that the token machine waits on. The synchronization done with the help of semaphores uses the same InfiniBand connection used to send KV-cache data. When processing a batch of prompts, each request is assigned a different semaphore since it may be routed to different token machines. We ship the KV-caches block-by-block in vLLM. To minimize the number of transfers, we also consider the contiguity of KV blocks as long as they use the same semaphore.
V-B Simulator setup
We build a simulator to explore cluster designs and evaluate Splitwise at scale. The simulator code is open source [20].
Figure 13: Overview of the design of the Splitwise simulator.
Figure 13 shows the design of our simulator. The simulator is event-driven and faithfully models the Splitwise machine pools, schedulers, machine-level memory and queues, and KV-cache transfer. We first profile the LLM on the target hardware with various input/output sizes  1. Based on the characterization profiles, we build a performance model. The simulator takes as input the request traces, SLOs, the performance model, and the configurations for cluster and scheduler  2. For our evaluation, we use the prompt and token size distributions from the production traces in Section III. We tune the Poisson arrival rate to increase and decrease the load (requests per second) for cluster sizing. The simulator provides the achieved metrics per request (TTFT, TBT, E2E), and the machine utilization levels  3. We cross-validated the performance model with hardware experiments to ensure accuracy; we also validated the simulator end-to-end using production load with over 50K iterations to ensure fidelity  4.
Performance model. We build a piece-wise linear performance model using performance profiles at various batch sizes, input sizes, output sizes, in the required parallelism configuration on A100 and H100 machines from Section III. We validate that our performance model has high accuracy; it incurs a mean absolute percentage error (MAPE) of less than 3% when evaluated with a 80:20 train:test dataset split.
Communication model. In our evaluation, KV-cache transfers cause inter-machine communication, whereas tensor parallelism only causes intra-machine communication. We model inter-machine communication overheads by benchmarking our KV-cache transfer implementation over Infiniband in Section VI-A.
SLOs. To determine the maximum throughput that can be supported by a given cluster design, we use P50, P90, and P99 SLOs for TTFT, TBT, and E2E latency metrics. Table VI shows our SLO definition using DGX-A100 as a reference. We require all nine SLOs to be met. SLOs on TTFT are slightly looser, since it has a much smaller impact on the E2E latency.
P50 P90 P99